OpenStudy (anonymous):

find the sum of the solutions of this equation 5x^2+3x+8=0

OpenStudy (anonymous):

Hey Jamae, to solve this there is a simple formula that works with every polynomial of the form \(ax^2+bx+c = 0\) \[ x_{1,2} = \frac{1}{2a} \left[ -b \pm\sqrt{b²-4ac} \right] \] For your equation this means the following: \[ \underbrace{5}_{a}x^2+\underbrace{3}_{b}x+\underbrace{8}_c = 0 \] if you plug your values in, you get the following: \[ x_{1,2} = \frac{1}{10}\left( - 3 \pm \sqrt{9-4\cdot5\cdot8} \right) = \frac{1}{10}\left( - 3 \pm \sqrt{9-4\cdot5\cdot8} \right) \] \[ = \frac{1}{10}\left( - 3 \pm \sqrt{-151} \right) \] \[ \Rightarrow x_1 =- \frac{1}{10}\mathbf{i}\left( \sqrt{151} - 3 \mathbf{i} \right), \quad x_1 = \frac{1}{10}\mathbf{i}\left( \sqrt{151}+ 3 \mathbf{i} \right) \]